OpenStudy (anonymous):

write product in its simples form -7x^3y^5*(-7y^3)

OpenStudy (anonymous):

49(x^3)(y^8) the exponents are added in multiplication

OpenStudy (fibonaccichick666):

not quite @Gretacig if \(-7x^{3y^5} \times -7y^3\) It is not quite the same

OpenStudy (fibonaccichick666):

@oburtono4 can you clarify the problem please

OpenStudy (anonymous):

you should definitely use more brackets. you can never have enough ^_^

OpenStudy (fibonaccichick666):

I quite =: [concur] ^.-

OpenStudy (anonymous):

SORRY, I MADE A MISTAKE. It should be good now: \[-7x ^{3y ^{5}}\times(-7y ^{3})=49x ^{3y ^{5}}\times y ^{3}\]